Bluebird Bakery produces bread sticks. A primary direct material is flour. Here are data regarding Bluebirds production activity for the month of May 100 pounds $1.20 per pound $1.35 per pound 5,000 pounds $5,700 favorable Standard quantity of flour per batclh Standard price per pound of flour Actual price paid for flour Actual quantity of flour purchased and used Flexible budget variance for flour What is the direct material quantity variance for flour? $5,250 unfavorable $4,950 favorable $6,450 favorable $7,650 unfavorable None of the above
Answer #1
Answer : $6450 favorable
=> DMQV = (Std. material allowed for actual production - actual usage) * Std. price
Std. material cost at actual output = actual cost + flexible budget variance = 5000*1.35 + 5700 = 12450
Std. allowed quantity material for actual output = 12450/1.2 = 10375 pounds
DMQV = (10375 - 5000)* 1.2 = $6450 F